OpenStudy (ronaldo7):

√y+5 + √y +13 = 4 Can someone help me with this too

OpenStudy (kendricklamar2014):

What have u done?

OpenStudy (ronaldo7):

what do you mean

OpenStudy (kendricklamar2014):

Have u tried to solve this questions or have u done any steps?

OpenStudy (ronaldo7):

oh yeah

OpenStudy (kendricklamar2014):

Can you post the steps?

OpenStudy (ronaldo7):

sure hold on

OpenStudy (ronaldo7):

I moved √y +5 to the other side then squared both sides

OpenStudy (ronaldo7):

like subtracteed it

OpenStudy (kendricklamar2014):

Well this is how I would solve it: \[√y+5 + √y +13 = 4 \] \[2√y+18=4\] \[2√y+18+−18=4+−18\] \[2√y=−14\] \[\frac{ 2√y }{ 2 } = \frac{ -14 }{ 2 }\] \[√y=−7\]

OpenStudy (kendricklamar2014):

\[y=(−7)^2\] \[y=49\]

OpenStudy (ronaldo7):

oh you must have misunderstood

OpenStudy (kendricklamar2014):

Answer: No Solution

OpenStudy (ronaldo7):

the radical is over the entire y +13 and the y +5

OpenStudy (ronaldo7):

Hello?